OpenStudy (anonymous):

Solve this limit algebraically: Lim as X approches 0 (X+sinX)/X

OpenStudy (anonymous):

2?

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{x+\sin(x)}{x}=\lim_{x \rightarrow 0}\frac{x}{x}+\lim_{x \rightarrow 0}\frac{\sin(x)}{x}=1+1=2\]

OpenStudy (anonymous):

\[\lim_{x\rightarrow 0}\frac{x}{x}+\frac{\sin(x)}{x}=1+1=2\]